# Efficient self-consistent learning of gate set Pauli noise

### v1

@@ I.2 Discussion and outlook
One promising practical application of our theory is in quantum error mitigation (QEM) [31]. There have been QEM protocols based on the Pauli noise models [14, 32, 33, 34] which work by first learning the Pauli noise channels and then either canceling the noise via quasi-probability sampling, or amplifying the noise and then extrapolating to the noiseless limit. Due to the existence of gauge, these protocols generally make assumptions that are not physically justified (e.g., “symmetry assumption” [14, 19] or perfect state preparation [34]) in order to determine the noisy Pauli channels. However, it is known that the issue of gauge ambiguity should not hinder self-consistent QEM if assisted by gate set tomography [35, 36], although the latter method is not easily scalable. Using our framework, by efficiently learning all the learnable degrees of freedom and anchoring at an arbitrary point of the gauge, one will be able to mitigate the Pauli noise without any additional assumptions. Such a self-consistent and efficient QEM protocol is actively being developed by some of the current authors.
@@
On the theory side, we believe that our framework and graph-theoretic analysis can be extended beyond the Pauli noise model. We expect the statement that the cycle space equals the learnable space to be held more generally. Note that a Pauli noise model has a linear parameterization (i.e., the log of the Pauli eigenvalues), and one can study the learnable and gauge degrees of freedom as linear subspaces of the parameter spaces. For more general noise models, one might need to study the tangent space of the parameters, which has been considered in the theory of the first-order gauge-invariant (FOGI) model [37, 38]. It is interesting to see if our method can be extended there to have a more quantitative understanding of the learnable and gauge degrees of freedom.
@@
Our framework resembles ACES (i.e., averaged circuit eigenvalue sampling), a Pauli noise learning protocol proposed by one of the authors [13] and further analyzed and implemented in Cisneros et al. [47] and Hockings et al. [20]. Similar to our setting, ACES models every different Clifford gate as having a gate-dependent Pauli noise channel (with an efficient noise ansatz such as fully local noise), and aims at learning every Pauli eigenvalue of every Pauli noise channel. ACES constructs linear equations of the form $\bm{b}=A\bm{x}$, where $\bm{x}$ are the logarithmic Pauli eigenvalues, $\bm{b}$ are the logarithmic circuit eigenvalues which can be experimentally measured, and $A$ is the design matrix determined by a set of Clifford circuits and Pauli measurements. The original ACES effectively fixes a gauge by ignoring state preparation noise, and is thus able to construct a full column-rank $A$ using random Clifford gates. Within ACES’s framework, our work basically says that if SPAM noise parameters are included in $\bm{x}$, there is a maximum possible rank of $A$ equal to the number of learnable degrees of freedom, and we provide an explicit recipe for constructing such an $A$. Moreover, the explicit construction can learn gate parameters to relative precision; see Theorem 4. It is interesting to extend the statistical analysis for ACES [20] to our setting. It is also interesting to study optimal learning and experimental design within our framework [19, 48]. Conversely, it is interesting to study the ultimate sample complexity limit for learning gate set Pauli noise models. Such bounds for learning Pauli channels in a black-box setting have been previously investigated [17, 49, 50, 51, 52].
The learning algorithm we proposed relies on the fact that the noise model of interest is realizable, i.e., there is no out-of-model error. It would be interesting to analyze the effect of such errors. Our current analysis and algorithms require all the noise to be Pauli channels with certain constraints, specifically, the locality constraints as discussed in Sec. VVI, and VII. In practice, one can fit the same data into models with different locality assumption, and there will be a tradeoff between model expressivity and trainability. One can also utilize techniques from structure learning to find the optimal (quasi-)local Pauli noise model (see e.g. [53]). It would be interesting to explore learning algorithms with those features within our framework.
@@ II Notations and Preliminaries
$\Omega$ Clique factor set of a Markov Random Field Sec. VI.1
$\Omega_{*}$ Subset of all maximal cliques from $\Omega$ Sec. VI.1
@@
Pauli group. We follow the notations of Ref. [9, 17]. Let ${\sf P}^{n}=\{I,X,Y,Z\}^{\otimes n}$ be the Pauli group modulo a global phase. ${\sf P}^{n}$ is an Abelian group isomorphic to $\mathbb{Z}^{2n}_{2}$. Specifically, given a $2n$-bit string $a=(a_{x},a_{z})=a_{x,1}\cdots a_{x,n}a_{z,1}\cdots a_{z,n}$, the corresponding Pauli operator is given by
@@
Pauli channel. An $n$-qubit Pauli channel $\Lambda$ has the following two equivalent forms
@@
Clifford group.- The $n$-qubit Clifford group ${\sf C}^{n}$ is the group of $n$-qubit unitary operators that preserve the Pauli group upon conjugation. Mathematically, $\forall C\in{\sf C}^{n}$, $\forall P_{a}\in{\sf P}^{n}$, $CP_{a}C^{\dagger}\in{\sf P}^{n}\times\{\pm\}$. We often denote the unitary channel of a Clifford gate using the same calligraphic letter, e.g., $\mathcal{C}(\rho)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=C\rho C^{\dagger}$. We also define, with slight abuse of notation, the mapping $\mathcal{C}:\mathbb{Z}_{2}^{2n}\mapsto\mathbb{Z}_{2}^{2n}$ via $\mathcal{C}(P_{a})\equiv CP_{a}C^{\dagger}=\pm P_{\mathcal{C}(a)}$. The Pauli-Liouville operator of a Clifford $\mathcal{C}$ has the following form,
@@
Miscellaneous. For a real linear space $X$, the dual space of it is denoted by $X^{\prime}$. For a linear map $\mathcal{Q}$, its dual map is denoted by $\mathcal{Q}^{{\sf T}}$. The image and kernel of a map $\mathcal{Q}$ are denoted by $\mathrm{Im}\mathcal{Q}$ and $\mathrm{Ker}\mathcal{Q}$, respectively. $\mathds{1}[X]$ is the indicator function that takes $1$ if the statement $X$ is true and 00 elsewise.
@@ III.1 Basic definitions
We comment that the learnable function we define here is also known as identifiable function in the literature of identifiability analysis [54]. Here, for consistency, we will stick to the term “learnable” which has been used in Ref. [15, 25].
@@ III.2 Learnability of complete noise model
Figure 2: The PTG for $n=2$, $\mathfrak{G}=\{\mathrm{CZ}\}$. Here we define $\mathcal{G}=\mathrm{CZ}$ for notational simplicity. The black edges corresponds to gate noise parameters, while the gray edges corresponds to SPAM noise parameters. Multiple edges are shown as one edge with multiple labels.
The following definition of algebraic graph theoretic terminology follows Ref. [15].
@@
A path is an alternating sequence of vertices and edges $z=(v_{i_{0}},e_{i_{1}},v_{i_{1}},e_{i_{2}},v_{i_{2}},...,v_{i_{q-1}},e_{i_{q}},v_{i_{q}})$ such that each edge satisfies $e_{i_{k}}=(v_{i_{k-1}},v_{i_{k}})$. A cycle is a closed path, which means $v_{i_{0}}=v_{i_{q}}$.111 In the literature of algebraic graph theory, people sometimes allow reverse edges (i.e., $e_{i_{k}}=(v_{i_{k}},v_{i_{k-1}})$) in the definition of a cycle and refer to our definition as an oriented cycle or a circuit. Since all graphs considered in this work are either strongly-connected or a union of strongly-connected components, and it has been shown that oriented cycles can linearly represent any cycle for such graphs [55], working with our current definition would not cause any inconsistency. A cycle is simple if it passes through each unique edge at most once. For a simple cycle $z$ in the PTG, we assign a vector $\bm{x}_{z}\in E$ as follows
@@
Remark: while $E=Z\oplus^{\perp}U$ is a graph-theoretic fact, we don’t know a priori whether $X=L\oplus T$ holds under our definition. This makes the proof different from [15].
@@ IV Algorithms for learning gate set Pauli noise
From this section, we focus on Clifford circuits as they are sufficient for learning Pauli noise. Also, instead of describing an experiment with its measurement probability distribution, we will talk about measuring certain Pauli observable. Since we assume noiseless single-qubit layers, measuring Pauli $P_{a}$ effectively picks up $\lambda^{M}_{\mathrm{pt}(a)}$ from the measurement noise channel. We thus define the effective noisy observable as $\tilde{P}_{a}\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\lambda^{M}_{\mathrm{pt}(a)}P_{a}$. Similarly, we use $\rho_{a}$ to denote the $+1$ eigenstate of $P_{a}$ obtained by applying an appropriate single-qubit Clifford layer to $\rho_{0}\equiv{|0\rangle\langle 0|}^{\otimes n}$.222 Although $\rho_{a}$ is not uniquely defined when $P_{a}$ contains identity, choosing any $\rho_{a}$ that satisfies this condition will make no observable difference. The noisy realization of $\rho_{a}$ is denoted as $\tilde{\rho}_{a}$ and satisfies $\operatorname{Tr}[P_{a}\tilde{\rho}_{a}]=\lambda^{S}_{\mathrm{pt}(a)}$. Finally, we use the tuple $({\rho_{a},\mathcal{C},P_{b}})$ to denote an experiment that (ideally) starts with $\rho_{a}$, applies the sequence of gates $\mathcal{C}$, and measures the observable $P_{b}$.
@@ V.2 Learnability of fully local noise model
Theorem 5(c) characterizes the learnability of fully local Pauli noise models. In words, the embedded gauge space $\mathcal{Q}(T_{R})$ is spanned exactly by those $n$ single-qubit depolarizing gauges. One can also write down the reduced gauge space as $T_{R}=\mathrm{span}\{\mathcal{Q}^{-1}(\mathfrak{d}_{s}):|s|=1\}$, where each $\mathcal{Q}^{-1}(\mathfrak{d}_{s})$ is well-defined333 Though $\mathcal{Q}$ is not invertible, it is injective and we know that $\forall|s|=1,~{}\mathfrak{d}_{s}\in{\mathrm{Im}{\mathcal{Q}}}$, thus $\mathcal{Q}^{-1}(\mathfrak{d}_{s})$ is uniquely defined. and can be efficiently calculated using the inverse transformation given in Lemma 5.4. The learnable space $L_{R}$ can be obtained by, e.g., computing the orthogonal complement of $T_{R}$ within $X_{R}$ using standard linear algebra procedures. We will come back to the construction of $L_{R}$ later.
Now we present the proof for Theorem 5(a), the direct part of 5(b), and 5(c). The converse part of 5(b) is slightly more involved and is postponed to App. C.2 for clarity.
@@ V.3 Efficient learning of fully local noise model
V.4 Relative-precision learning algorithms for 2-qubit gates
Now we consider the second type of algorithms that learn gate noise parameters to relative precision. The key is to construct a reduced cycle basis for the gate noise parameter space $L_{R}^{G}=\mathcal{Q}^{\sf T}(L\cap X^{G})$. Note that, one might think
We find this problem to be hard in general. However, if every $\mathcal{G}\in\mathfrak{G}$ is supported non-trivially on at most $2$ qubits, we do have an efficient algorithm to solve this problem. Since many existing quantum computing platform uses two-qubit entangling gates such as CZ as a primitive, we expect our method to be practically useful.
Algorithm 3  [Finding a basis for $\widetilde{L_{R}^{G}}$].
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Repeat step 2 and 3 until no $\bm{c}$ can be found, return $C$.
Lemma 5.8.
Algorithm 3 outputs a basis for $\widetilde{L_{R}^{G}}$.
This is a standard linear algebraic procedure. For completeness, a proof is included in App. C.3.
Definition 5.9.
For $\bm{r}\in X_{R}^{G}$, $\partial(\bm{r})$ is called its boundary.
Lemma 5.10.
Proof.
$\Leftrightarrow$ $\quad\partial(\bm{r})=\bm{0}.$
In the first stage, we look for $\bm{c}=\sum_{a\in\sf P^{2}\backslash\{I_{2}\},\mathcal{G}\in\mathfrak{G}}r_{a}^{\mathcal{G}}\bm{\vartheta}_{a}^{\mathcal{G}}$ with the property that all non-zero $r_{a}^{\mathcal{G}}$ share the same $\mathrm{supp}(\mathcal{G})$. That is, we group the base vectors $\bm{\vartheta}_{a}^{\mathcal{G}}$ by $\mathrm{supp}(\mathcal{G})$ and consider them separately. Fix a common support $s$, the task simplifies substantially since Lemma 5.10 simplifies into a requirement on $2$ dimensions (the common support, which is of size $2$). Then we may perform the analysis on the pattern transfer subgraph (the union of the pattern transfer subgraphs of $\hat{\mathcal{G}}$ for all $\mathcal{G}\in\mathfrak{G}$ with $\mathrm{supp}(\mathcal{G})=s$). In the following we consider the possible shapes of the pattern transfer subgraph. Since self loops can be indentified and picked as reduced cycle basis easily at this stage, we do not need to explicitly consider them.
Lemma 5.11.
Proof.
From Figure 4 one can easily see that the union of pattern transfer subgraphs still has only $3$ possible shapes. The important message of Lemma 5.11 is that for any edge in the pattern transfer subgraph, there is another edge of the reverse direction.
$\Downarrow$
(a) There exists two type-1 boundaries with the same support.
{NiceTabular}
$\Downarrow$
[…]
(b) Chasing procedure at both end terminates with a type-1 boundary.
{NiceTabular}
$\Downarrow$
[…]
(c) A chasing procedure ends with a loop.
Now the algorithm proceeds to the second stage, where we find reduced cycles across different supports. This stage is illustrated in Figure 5. According to Lemma 5.10 we associate the base vectors in $A$ with their boundaries. Thanks to Lemma 5.11, there are only two types of boundaries: Type 1 boundary takes ±$\pm 1$ at one entry and 00 at other places, corresponding to edges (10,11) or (01,11) in Fig. 4. Type $2$ boundary takes $1$ at one entry, $-1$ at another entry, and 00 at other places, corresponding to edge (10,01) in Fig. 4. Moreover, if we consider the edge of reverse direction, the corresponding boundary is negated. In the following we will find some candidate $\bm{c}=\sum\mathrm{sgn}_{i}\bm{\vartheta}_{i}$ and then prove that it is a reduced cycle when all the signs are positive. When some signs are negative, we revert some edges to make the signs positive. This is actually a simplification and hides the following fact: though $\bm{c}$ itself is not a reduced cycle, $\tilde{\bm{c}}=\bm{c}+\sum_{\mathrm{sgn}_{i}=-1}(\bm{\vartheta}_{i}+\bm{\vartheta}_{i}^{\prime})$ ($\bm{c}$ with edge directions reverted) is a reduced cycle. Moreover, since the $\bm{\vartheta}_{i}+\bm{\vartheta}_{i}^{\prime}$s are already in $C$, putting $\bm{c}$ or $\tilde{\bm{c}}$ into $C$ is equivalent in terms of making $C$ a basis for $\widetilde{L_{R}^{G}}$.
According to the proof of Lemma 5.8, if there are remaining degree of freedom in $\widetilde{L_{R}^{G}}$, there should exist a non-zero vector $\bm{c}^{\prime}\in\widetilde{L_{R}^{G}}\cap\mathrm{supp}(A)$. Then $\partial(\bm{c}^{\prime})=\bm{0}$. We can find such $\bm{c}^{\prime}=\sum_{i}r_{a_{i}}^{\mathcal{G}_{i}}\bm{\vartheta}_{a_{i}}^{\mathcal{G}_{i}}$ (the summation is only over non-zero terms) via simple linear algebra on the boundaries of the edges in $A$. The $r_{a_{i}}^{\mathcal{G}_{i}}$ found is rational and thus can be made integral by multiplying them by the least common multiple of the denominators.
@@ VI.1 Basic definitions
Given an undirected graph with $n$ vertices labeled by $[n]$, called a factor graph, we use $\Omega$ to denote the set of all cliques, i.e., subsets of vertices where each pair of vertices are connected by an edge. Equivalently, $\Omega$ is a subset of $2^{[n]}$ that satisfies $\forall\nu\in\Omega$, $\forall\varnothing\subsetneq\mu\subseteq\nu$, $\mu\in\Omega$. We call such $\Omega$ a clique factor set. We also use $\Omega_{*}$ to denote the set of maximal cliques within $\Omega$, i.e., those that do not belong to any other cliques. As an example, $\Omega=\{\{1\},\{2\},\{3\},\{1,2\},\{2,3\}\}$ is a valid clique factor set, with the corresponding $\Omega_{*}=\{\{1,2\},\{2,3\}\}$. Obviously, $\Omega$, $\Omega_{*}$ uniquely determines each other.
@@
for some $\{r_{b}\in\mathbb{R}:b\sim\Omega,b\neq I_{n}\}$ and a clique factor set $\Omega$.
@@
A few remarks regarding this definition: First, all $n$-qubit Pauli channels (with all Pauli eigenvalues being positive) are $2^{[n]}\backslash\varnothing$-local. Note that we do not require $r_{b}$ to be non-negative. Second, this definition basically says that the eigenvalues $\{\lambda_{a}\}_{a}$ factorize according to a Markov Random Field (MRF) [65] specified by the clique decomposition $\Omega$. This is opposed to [9] where Pauli channels with error rates $\{p_{a}\}_{a}$ factorized according to an MRF are considered. It is unclear whether these two models are equivalent, even approximately. Alternatively, our definition of $\Omega$-local is equivalent to requiring that $\Lambda$ can be expressed as a concatenation of Pauli diagonal maps444 Note that, we do not require those Pauli diagonal maps to be completely positive. Similarly, when we say our quasi-local model is equivalent to the sparse Pauli-Lindblad model, we have not required the positivity of the model parameters. acting on each subsystem from $\Omega$, as proven in [16] (In fact, this is the picture we will use in most figures shown in this section). One can also show that the sparse Pauli-Lindblad model used in [14, 33, 66] is equivalent to our definition. Finally, throughout this paper we assume $\Omega$ is known a priori, from, e.g., knowledge about the device topology, leaving the topic of structure learning (see e.g. [53]) for future study. We provide more details about the quasi-local Pauli noise models in Appendix D.

### v2

@@ I.3 Future directions
Quantum error mitigation. One promising practical application of our theory is in quantum error mitigation (QEM) [40]. There have been QEM protocols based on the Pauli noise models [15, 41, 42, 43] which work by first learning the Pauli noise channels and then either canceling the noise via quasi-probability sampling, or amplifying the noise and then extrapolating to the noiseless limit. Due to the existence of gauge, these protocols generally make assumptions that are not physically justified (e.g., “symmetry assumption” [15, 19] or perfect state preparation [43]) in order to determine the noisy Pauli channels. However, it is known that the issue of gauge ambiguity should not hinder self-consistent QEM if assisted by gate set tomography [44, 45], although the latter method is not easily scalable. Using our framework, by efficiently learning all the learnable degrees of freedom and anchoring at an arbitrary point of the gauge, one will be able to mitigate the Pauli noise without any additional assumptions. Such a self-consistent and efficient QEM protocol is actively being developed by some of the current authors [46].
@@
Beyond Pauli noise model. We believe that our framework and graph-theoretic analysis can be extended beyond the Pauli noise model. We expect the statement that the cycle space equals the learnable space to be held more generally. Note that a Pauli noise model has a linear parameterization (i.e., the log of the Pauli eigenvalues), and one can study the learnable and gauge degrees of freedom as linear subspaces of the parameter spaces. For more general noise models, one might need to study the tangent space of the parameters, which has been considered in the theory of the first-order gauge-invariant (FOGI) model [47, 48]. It is interesting to see if our method can be extended there to have a more quantitative understanding of the learnable and gauge degrees of freedom.
@@
Beyond realizable setting. The learning algorithm we proposed relies on the fact that the noise model of interest is realizable, i.e., there is no out-of-model error. It would be interesting to analyze the effect of such errors. Our current analysis and algorithms require all the noise to be Pauli channels with certain constraints, specifically, the locality constraints as discussed in Sec. VVI, and VII. In practice, one can fit the same data into models with different locality assumption, and there will be a tradeoff between model expressivity and trainability. One can also utilize techniques from structure learning to find the optimal (quasi-)local Pauli noise model (see e.g. [57]). It would be interesting to explore learning algorithms with those features within our framework.
@@ II Notations and Preliminaries
$\Omega$ Factor set of a Markov Random Field Sec. VI.1
$\Omega_{*}$ Subset of all maximal factors from $\Omega$ Sec. VI.1
@@
Pauli group. We follow the notations of [9, 17]. Let ${\sf P}^{n}=\{I,X,Y,Z\}^{\otimes n}$ be the Pauli group modulo a global phase. ${\sf P}^{n}$ is an Abelian group isomorphic to $\mathbb{Z}^{2n}_{2}$. Specifically, given a $2n$-bit string $a=(a_{x},a_{z})=a_{x,1}\cdots a_{x,n}a_{z,1}\cdots a_{z,n}$, the corresponding Pauli operator is given by
@@
Pauli channel.
An $n$-qubit Pauli channel $\Lambda$ has the following two equivalent forms
@@
Clifford group. The $n$-qubit Clifford group ${\sf C}^{n}$ is the group of $n$-qubit unitary operators that preserve the Pauli group upon conjugation. Mathematically, $\forall C\in{\sf C}^{n}$, $\forall P_{a}\in{\sf P}^{n}$, $CP_{a}C^{\dagger}\in{\sf P}^{n}\times\{\pm\}$. We often denote the unitary channel of a Clifford gate using the same calligraphic letter, e.g., $\mathcal{C}(\rho)\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=C\rho C^{\dagger}$. We also define, with slight abuse of notation, the mapping $\mathcal{C}:\mathbb{Z}_{2}^{2n}\mapsto\mathbb{Z}_{2}^{2n}$ via $\mathcal{C}(P_{a})\equiv CP_{a}C^{\dagger}=\pm P_{\mathcal{C}(a)}$. The Pauli-Liouville operator of a Clifford $\mathcal{C}$ has the following form,
@@
Miscellaneous. For a real linear space $X$, the dual space of it is denoted by $X^{\prime}$. For a linear map $\mathcal{Q}$, its dual map is denoted by $\mathcal{Q}^{{\sf T}}$. The image and kernel of a map $\mathcal{Q}$ are denoted by $\mathrm{Im}\mathcal{Q}$ and $\mathrm{Ker}\mathcal{Q}$, respectively. $\mathds{1}[X]$ is the indicator function that takes $1$ if the statement $X$ is true and 00 elsewise. All $\log$ represents natural logarithm.
@@ III.1 Basic definitions
We comment that the learnable function we define here is also known as identifiable function in the literature of identifiability analysis [58]. Here, for consistency, we will stick to the term “learnable” which has been used in [16, 26].
@@ III.2 Learnability of complete noise model
The following definition of algebraic graph theoretic terminology follows [16].
@@
A path is an alternating sequence of vertices and edges $z=(v_{i_{0}},e_{i_{1}},v_{i_{1}},e_{i_{2}},v_{i_{2}},...,v_{i_{q-1}},e_{i_{q}},v_{i_{q}})$ such that each edge satisfies $e_{i_{k}}=(v_{i_{k-1}},v_{i_{k}})$. A cycle is a closed path, which means $v_{i_{0}}=v_{i_{q}}$.333 In the literature of algebraic graph theory, people sometimes allow reverse edges (i.e., $e_{i_{k}}=(v_{i_{k}},v_{i_{k-1}})$) in the definition of a cycle and refer to our definition as an oriented cycle or a circuit. Since all graphs considered in this work are either strongly-connected or a union of strongly-connected components, and it has been shown that oriented cycles can linearly represent any cycle for such graphs [28], working with our current definition would not cause any inconsistency. A cycle is simple if it passes through each unique edge at most once. For a simple cycle $z$ in the PTG, we assign a vector $\bm{x}_{z}\in E$ as follows
@@
Remark: while $E=Z\oplus^{\perp}U$ is a graph-theoretic fact, we don’t know a priori whether $X=L\oplus T$ holds under our definition.
@@ IV Algorithms for learning gate set Pauli noise
From this section, we focus on Clifford circuits as they are sufficient for learning Pauli noise. Also, instead of describing an experiment with its measurement probability distribution, we will talk about expectation values of measuring certain Pauli observable. Since we assume noiseless single-qubit layers, measuring Pauli $P_{a}$ effectively picks up $\lambda^{M}_{\mathrm{pt}(a)}$ from the measurement noise channel. We thus define the effective noisy observable as $\tilde{P}_{a}\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\lambda^{M}_{\mathrm{pt}(a)}P_{a}$. Similarly, we use $\rho_{a}$ to denote the $+1$ eigenstate of $P_{a}$ obtained by applying an appropriate single-qubit Clifford layer to $\rho_{0}\equiv{|0\rangle\langle 0|}^{\otimes n}$.444 Although $\rho_{a}$ is not uniquely defined when $P_{a}$ contains identity, choosing any $\rho_{a}$ that satisfies this condition will make no observable difference. The noisy realization of $\rho_{a}$ is denoted as $\tilde{\rho}_{a}$ and satisfies $\operatorname{Tr}[P_{a}\tilde{\rho}_{a}]=\lambda^{S}_{\mathrm{pt}(a)}$. Finally, we use the tuple $({\rho_{a},\mathcal{C},P_{b}})$ to denote an experiment that (ideally) starts with $\rho_{a}$, applies the sequence of gates $\mathcal{C}$, and measures the observable $P_{b}$.
@@ V.2 Learnability of fully local noise model
Theorem 10(c) characterizes the learnability of fully local Pauli noise models. In words, the embedded gauge space $\mathcal{Q}(T_{R})$ is spanned exactly by those $n$ single-qubit depolarizing gauges. One can also write down the reduced gauge space as $T_{R}=\mathrm{span}\{\mathcal{Q}^{-1}(\mathfrak{d}_{s}):|s|=1\}$, where each $\mathcal{Q}^{-1}(\mathfrak{d}_{s})$ is well-defined555 Though $\mathcal{Q}$ is not invertible, it is injective and we know that $\forall|s|=1,~{}\mathfrak{d}_{s}\in{\mathrm{Im}{\mathcal{Q}}}$, thus $\mathcal{Q}^{-1}(\mathfrak{d}_{s})$ is uniquely defined. and can be efficiently calculated using the inverse transformation given in Lemma 5.4. The learnable space $L_{R}$ can be obtained by, e.g., computing the orthogonal complement of $T_{R}$ within $X_{R}$ using standard linear algebra procedures. We will come back to the construction of $L_{R}$ later.
Now we present the proof for Theorem 10(a), the direct part of 10(b), and 10(c). The converse part of 10(b) is slightly more involved and is postponed to App. C.2 for clarity.
@@ V.3 Efficient learning of fully local noise model
To learn the gate noise parameters to relative precision, one needs to construct a reduced cycle basis for the gate noise parameter space $L_{R}^{G}=\mathcal{Q}^{\sf T}(L\cap X^{G})$. We find this problem hard in general. However, if every $\mathcal{G}\in\mathfrak{G}$ is supported non-trivially on at most 2 qubits, we do have an efficient algorithms to construct such a basis. See App D.
@@ VI.1 Basic definitions
Let $\Omega$ be a subset of $2^{[n]}$ that satisfies $\forall\nu\in\Omega$, $\forall\varnothing\subsetneq\mu\subseteq\nu$, $\mu\in\Omega$. We call such $\Omega$ a factor set. We also use $\Omega_{*}$ to denote the set of maximal factors within $\Omega$, i.e., those that do not belong to any other factor. As an example, $\Omega=\{\{1\},\{2\},\{3\},\{1,2\},\{2,3\}\}$ is a valid factor set, with the corresponding $\Omega_{*}=\{\{1,2\},\{2,3\}\}$. Obviously, $\Omega$, $\Omega_{*}$ uniquely determines each other.
@@
for some $\{r_{b}\in\mathbb{R}:b\sim\Omega,b\neq I_{n}\}$ and a factor set $\Omega$.
@@
A few remarks regarding this definition: First, all $n$-qubit Pauli channels (with all Pauli eigenvalues being positive) are $2^{[n]}\backslash\varnothing$-local. Note that we do not require $r_{b}$ to be non-negative. Second, this definition basically says that the eigenvalues $\{\lambda_{a}\}_{a}$ factorize according to a Markov Random Field (MRF) [68] specified by the factor decomposition $\Omega$. This is opposed to [9] where Pauli channels with error rates $\{p_{a}\}_{a}$ factorized according to an MRF are considered. It is unclear whether these two models are equivalent, even approximately. Alternatively, our definition of $\Omega$-local is equivalent to requiring that $\Lambda$ can be expressed as a concatenation of Pauli diagonal maps666 Note that, we do not require those Pauli diagonal maps to be completely positive. Similarly, when we say our quasi-local model is equivalent to the sparse Pauli-Lindblad model, we have not required the positivity of the model parameters. acting on each subsystem from $\Omega$, as proven in [14] (In fact, this is the picture we will use in most figures shown in this section). One can also show that the sparse Pauli-Lindblad model used in [15, 42, 69] is equivalent to our definition. Finally, throughout this paper we assume $\Omega$ is known a priori, from, e.g., knowledge about the device topology, leaving the topic of structure learning (see e.g. [57]) for future study. We provide more details about the quasi-local Pauli noise models in Appendix E.